# HAAP: Vision-context Hierarchical Attention Autoregressive with Adaptive Permutation for Scene Text Recognition

### v1

@@ III-C Implicit Permutation Neurons (IPN)
Considering that STR traditionally relies on left-to-right $P_{left2right}$ or right-to-left $P_{right2left}$ unidirectional sequence modeling, we propose IPN to improve the correlation between inputs and outputs through multidirectional sequence modeling. Essentially, multidirectional sequence modeling considers internal correlations while considering bi-directionality to avoid the effects of word syntactic structures or external disturbances such as occlusions and distortions. For example, to predict the ”e” in the word ”clearance”, we consider ”cl_” and ”ecnara_”. ” in the context of ”cl_ar”. IPN provides an attention mask during the attention operation to generate dependencies between the input context and the output without actually replacing the text labels. Table. I illustrates three examples of the mask.
@@
Formally, first, given the $T$-length text labels $y=[y^{1},y^{2},\cdots,y^{T}]$, decompose the likelihood $[1,2,\cdots,T]$ according to the canonical order using the chain rule to build a left-to-right permutation of $P_{left2right}$ to obtain the model inference as:
@@
where $M_{w}$ is the weighted transformation matrix. This process helps the model to parse the representation of the structure of the input data. Even though there is no explicit nonlinear activation function throughout the process, the model is still able to capture the nonlinear relationships in the data because of the learning of the weights. Fourth, the transformation matrix is inversely mapped to the sequence of scores $P_{score}$ via $P_{value}$, as:
@@
Next, $P_{score}$ is normalized from largest to smallest to generate the actual ranking order. Immediately, we invert the left-to-right permutations $[1,2,.,T]$ and invert the adaptive permutation pairs to obtain twice the amount of information.
@@ III-D Cross-modal Hierarchical Attention mechanism (CHA)
The standard Transformer parametric AR model is inefficient for decoding in multidirectional sequence modeling since the feature distribution predicted by the hidden state is independent of the target location [41]. Let us assume that we parameterize the next token distribution as:
@@
where $h_{\theta}(x_{r<t})$ denotes the hidden representation of $x_{r<t}$ produced by the shared transformer network after proper masking. We use $rt$ and $r<t$ to denote the $t$-th and the first $t-1$ elements of the permutation set $r$. Since the same model parameter $\theta$ is shared between all decomposition sequences during training, $rt$ is expected to see that every possible element in the sequence is not equal to $x_{t}$. Note that the representation $h_{\theta}(x_{z<t})$ does not depend on the position it will predict, i.e. the value of $rt$. Two different target positions share the same model predictions. However, the ground truth distributions for the two positions are certainly different.
To avoid this problem, we perform context and location encoding based on a two-stream attention mechanism to provide complete contextual information. We use two sets of hidden representations i.e., from content and query representations. Meanwhile, considering that context and location need to interact with visual features cross-modally for alignment, we adopt a hierarchical coupling approach, i.e., CHA, to consider the context-location-image association. In the first stage, context-visual attention is established after self-attentive encoding of sorted contexts, as:
$Attn_{cv}=p+dropout[MHA(Attn_{c},z)],$ (12)
@@
where $c\in\mathbb{R}^{(T+1)\times d_{model}}$ is the contextual embedding with positional information and $M\in\mathbb{R}^{(T+1)\times(T+1)}$ is the attention mask. The total sequence length is increased to $T+1$ by using special separator markers.
Next, MHA is used to establish positional-contextual-visual attention, as:
$Attn_{f}=Attn_{p}+dropout[MHA(Attn_{p},z)],$ (14)
@@
$Attn_{p}=p+dropout[MHA(p,Attn_{c}v,M)].$ (15)
The strategy of hierarchical feature processing prompts sequence interactions while utilizing image-guided context to assist the model in capturing dependencies between locations. The location-context-image interaction capability is boosted since adaptive permutation guides location interconnections. The robustness of the model is improved because the uniform gradient flow allows them to interact with each other during training. Finally, the output logits $y^{\prime}$ are obtained as:
@@
where $S$ is the size of the character set used for training. Additional characters are associated with the [E] tag (the end of the tagging sequence). The complete training loss for the optimization phase is the average of the $K$ cross-entropy loss $\mathcal{L}_{ce}$ with attention mask, as:
@@ IV-A Dataset
This experiment uses both synthetic training datasets MJSynth (MJ) [27] and SynthText (ST) [42] as well as real data for training including nine real-world datasets (COCO-Text (COCO) [43], RCTW17 [44], Uber-Text (Uber) [45], ArT [46], LSVT [47], MLT19 [48], ReCTS [49], TextOCR [50] and OpenVINO [51]). We use IIIT 5k-word (IIIT5k) [52], CUTE80 (CUTE) [53], Street View Text (SVT) [54], SVT-Perspective (SVTP) [55], ICDAR 2013 (IC13) [56] and ICDAR 2015 (IC15) [57], COCO, Uber, ArT for testing. The dataset contains challenging texts such as multi-orientation, curvature, perspective blur, low resolution, distortion, and occlusion.
@@
Synthetic datasets MJ consists of 8.9 million photorealistic text images composed of artificial datasets through over 90k English dictionaries. The composition of MJ consists of three parts background, foreground, and optional shadow/border respectively. It uses 1400 different fonts. Additionally, the font word spacing, thickness, underlining, and other attributes of MJ are different. MJ also utilizes different background effects, border/shadow rendering, base shading, projection distortion, natural image blending, and noise. ST consists of 8 million word text images synthesized by blending text over natural images. It uses scene geometry, textures, and surface normals to naturally blend and distort text renderings on the surfaces of objects in the image. Similar to MJ, ST uses randomized fonts for its text. The text image is cropped from the natural image in which the synthesized text is embedded.
@@
Real-world datasets The COCO dataset has a total of 73,127 training and 9.8k test images containing non-text, legible text, illegible text, and occluded text images. Each image contains at least one instance of legible text. RCTW contains 12,263 annotated images of the large-scale Chinese field dataset. The Uber dataset has a total of 127,920 training 60,000 test images collected from Bing Maps Streetside. It contains house numbers and vertical and rotated text on sign boards. The ArT dataset was created to recognize arbitrarily shaped text containing 32,028 training and 26,000 testing images of perspective, rotated, or curved arbitrarily shaped text. SVT is a large-scale street scene text dataset collected from streets in China totaling 41,439 text images of arbitrarily shaped natural scenes. MLT19 was created to recognize multilingual text. It has a total of 56,727 training images consisting of seven languages: Arabic, Latin, Chinese, Japanese, Korean, Bengali and Hindi. The ReCTS dataset contains 26,432 irregular texts arranged in various layouts or written in unique fonts. TextOCR and OpenVINO are large datasets with very diverse images containing 818,087 and 2,071,541 images respectively. They have complex scenes with multiple objects and text of different resolutions, orientations, and qualities.
@@
The IIIT5k dataset is a collection of 5000 natural scene and digitally born text images crawled from Google image search containing 2,000 images for evaluation and 3,000 images for testing. It has low resolution, multiple font styles, light and dark transformations, and projection distortion. CUTE contains 288 cropped images for the curved text collection. The images were captured by digital cameras or collected from the Internet. SVT is a collection of street view text from Google Street View. It contains 257 images for evaluation and 647 images for testing. Further, the SVTP dataset is a more complex collection of images containing numerous perspective texts totaling 645 images for testing. IC13 and IC15 were created for the ICDAR Robust Reading Contest. IC13 contains 848 images for training and 1,015 images for evaluation. IC15 contains 4,468 images for training and 2,077 images for evaluation. Many of the samples contain perspective, and blurred text.
@@ IV-B Implementation Details
The entire training and testing is implemented in two RTX 3090 GPUs with 24GB of RAM based on the Pytorch library [59]. The training is performed with a batch size of 1024 for 63,630 iterations i.e. 20 epochs on a real dataset of 3,257,585 samples and 4 epochs on a synthetic dataset of 16.89M samples. Following previous work [12], the Adam [60] optimizer is used along with the 1cycle [61] learning rate scheduler with an initial learning rate of $57e-$ 5. We set a maximum label length of $T=25$ and used a character set of size $S=94$ containing mixed-case alphanumeric characters and punctuation. The image is enhanced, resized, and normalized. First, the image enhancement is done using a three-layer RandAugment [62] operation where sharpening is tuned to Gaussian Blur and Poisson Noise. Second, the image is unconditionally resized to $128\times 32$ pixels, and patches of $8\times 4$ size are used. For the model testing phase, we use a 36-character set i.e. containing numbers and 26 letters. We use word accuracy as the main evaluation metric i.e., the prediction is considered correct if and only if the characters match at all positions.
Remark The highlighted red and red lines in the visualization results indicate error recognition and omissions, respectively. The best performance result is shown in bold font.
@@ IV-C Comparison with State-of-the-Arts
We compare HAAP with recent methods on 6 public benchmarks, and the quantitative results are listed in Table. II. We can see that HAAP outperforms the benchmark Parseq [12], realizing SOTA performance. The recognition performance in IC15 (incident scene text), SVTP (perspective scene text), and CUTE80 (curved text line image) are improved by $1.1\%$, $1.7\%$, and $1\%$ respectively. Our model achieves $99.3\%$ recognition accuracy on the IIIT5K (distorted, low-resolution scene text) dataset. This is an improvement of $2.6\%$ compared to the latest method, HVSI [35]. This verifies the excellent performance of HAAP on irregular text datasets, which contain a large number of low-quality images such as noisy and blurred images.
Besides validating the model on small-scale public benchmarks, we evaluate HAAP on three larger and more challenging recent benchmarks consisting of irregular text with different shapes, low-resolution images, rotations, and occlusions. The results are shown in Table. III. It achieves $80.6\%$, $85.5\%$, and $85.8\%$ outperforms the previous methods in terms of accuracy, respectively. Representative visible results are shown in Fig. 6, which demonstrates that our method is sufficiently robust to occlusion and text direction variability. In detail, we find that HAAP and recent methods are effective in solving complex background problems caused by perspective shifts or uneven lighting. However, artistic lettering and occlusion can lead to errors in previous work. The independence of heterogeneous data leads to ambiguous judgments by the recognizer, which can be effectively avoided by HAAP. Furthermore, HAAP can protect against additional distortions and ambiguities. For example, although the second half of ”CORNER” in Fig. 6 is ambiguous, our model can still reason effectively. HAAP also recognizes that the ”R” in ”SHARP” in Fig. 6 relies on AR reasoning with adaptive permutations.
@@
In addition, Fig. 5 and Table. IV shows the trade-off between accuracy and cost (parameters, FLOPS, and latency). For concise representation, we use $Avg$ to denote the average accuracy, which is computed as a weighted average based on the number of datasets. HAAP achieves the highest average word accuracy and exhibits competitive cost quality in all three metrics. Compared to PARSeq, HAAP uses far fewer parameters and FLOPs. In terms of latency, HAAP achieves $Avg-6$ of 95.0% with 15.2 ms/image, which outperforms previous methods.
TABLE IV: Comparison with other methods in terms of complexity and efficiency. $AVG-6$ represents the average accuracy in the six benchmarks.
@@
Ours 23.8 3.2 15.2 95.0
IV-D Ablation study
To evaluate the effectiveness of the proposed HAAP with IPN and CHA, we conduct ablation studies on nine benchmark datasets. In these experiments, all models are trained using real datasets with consistent settings of their model hyper-parameters.
The effectiveness of IPN To study the impact of adaptive permutations generated by IPN, we compare the sequence modeling strategies using left-to-right, bi-directional, and PLM. From Table. V, we find that the model using the bidirectional can achieve about $3.4\%$ improvement in $Avg-9$ compared to the left-to-right strategy. This proves that the double amount of information brought by bidirectional sequence modeling can effectively improve the correlation ability between sequence contexts. The introduction of PLM allows the model to achieve $90.12\%$ performance in terms of $Avg-9$ along with an additional 4 sets of random permutations. Immediately, the model with IPN achieves $0.28\%$ performance gain in terms of $Avg-9$ compared to the model with PLM. Meanwhile, as shown in Fig. 7, the model with IPN outperforms the introduction of PLM in terms of stability. This indicates that adaptive multi-directional sequence modeling can avoid the effects caused by the grammatical structure of words or by external disturbances. It also prevents uncertainty with human interference. The Comparison of qualitative results is shown in Fig. 8.
TABLE V: Ablation study of the IPN.$Avg-9$ represents the average accuracy in nine datasets. $P_{left2right}$ and $P_{right2left}$ represent the left-to-right and right-to-left permutation.
@@ IV-D Ablation study
Figure 7: Comparison of Loss (Left) and Valuation (Right) curves with IPN or PLM.
The effectiveness of CHA Considering that the unified gradient flow allows the modules to interact with each other, we evaluate CHA and MHA while evaluating their impact in different sequencing sequences. As shown in Table. VI, we find that the model with joint PLM and MHA (PLM+MHA) requires additional IR to improve the generalization. The PLM+CHA model achieves $90.15\%$ performance in terms of $Avg$ without additional IR, outperforming the performance of the PLM+MHA model with 3 number of refinement iterations ($90.12\%$). In addition, compared to the PLM+MHA model, the IPN+CHA model achieves a performance gain of $0.59\%$ without the need for refinement iterations. Besides, we find that the contribution of IR is minimal by introducing CHA in any sequence modeling (PLM or IPN). Because of the uniform gradient flow, the IPN+CHA model achieves the performance of $90.40\%$ and gets a performance gain of $0.25\%$ compared to the PLM+CHA model. This suggests that CHA can enhance the model generalization by providing effective positive feedback to the adaptive permutation.

### v2

@@ III-C Implicit Permutation Neurons (IPN)
Considering that STR traditionally relies on left-to-right $P_{left2right}$ or right-to-left $P_{right2left}$ unidirectional sequence modeling, we propose IPN to enhance the correlation between inputs and outputs through multidirectional sequence modeling. Essentially, multidirectional sequence modeling considers internal correlations of context while keeping attention to the bi-directional information flow to capture the word language rules and avoid the effort of visual interference (i.g, occlusion and distortion). For example, to predict the first “e” in the word “clearance,” we consider “cl_,” “ecnara_,” and the context of “cl_ar”. IPN provides an attention mask during the attention operation to generate dependencies between the input context and the output without actually replacing the text labels. Table. I illustrates three examples of the mask.
@@
Formally, first, given the $T$-length text labels $y=[y^{1},y^{2},\cdots,y^{T}]$, decompose the likelihood $1,2,\cdots,T]$ according to the canonical order using the chain rule to build a left-to-right permutation of $P_{left2right}$ to obtain the model inference as:
@@
where $M_{w}$ is the weighted transformation matrix. This process helps the model capture the input representation. Even though there is no explicit nonlinear activation function throughout the process, the model can still capture the nonlinear relationships in the data because of the learning of the weights. The transformation matrix is then mapped to the sequence of scores $P_{score}$ via $P_{value}$, as:
@@
Next, $P_{score}$ is normalized from largest to smallest to generate the actual ranking order. Immediately, we invert the left-to-right $[1,2,.,T]$ and adaptive permutation to obtain twice the amount of information.
@@ III-D Cross-modal Hierarchical Attention mechanism (CHA)
In multi-directional sequence modeling, the standard Transformer-based autoregressive model is inefficient since the feature distribution predicted by the hidden state is independent of the target position [54]. Let us assume that we parameterize the next token distribution as:
@@
where $h_{\theta}(x_{r<t})$ denotes the hidden representation of $x_{r<t}$ produced by the shared transformer network after proper masking. We use $x_{rt}$ and $x_{r<t}$ to denote the $t$-th and the first t1t-1 elements of the permutation set $r$. Since the same model parameter $\theta$ is shared between all decomposition sequences during training, $x_{rt}$ is expected to see every possible element in the sequence. Note that the position token $x_{rt}$ to be predicted only depends on the representation $h_{\theta}(x_{r<t})$. If only using the initial contextual semantics learned from the multi-directional sequence modeling without considering the current input visual information, it is not possible to update the representation $h_{\theta}(x_{r<t})$ with global information, increasing the probability of error accumulation. For this, researchers employ IR to improve the representation of $h_{\theta}(x_{r<t})$ [11, 4]. The IR uses the cloze mask to enhance the understanding of $h_{\theta}(x_{r<t})$ for the context because the cloze mask is able to provide the rest of the information other than the current position.
To avoid this problem, we introduce CHA to hierarchically encode context and position by allowing context, position, and image to interact with each other through a two-stream attention mechanism. We use two sets of hidden representations, i.e., from contextual semantic and query representations. In the first stage, context-image attention is established, producing the context semantic representation, which encodes both the context and $x_{rt}$ itself tokens guided by visual information, as:
$Attn_{cv}=Attn_{c}+dropout[MHA(Attn_{c},z_{l})],$ (9)
@@
where $c\in\mathbb{R}^{(T+1)\times d_{model}}$ is the context embeddings with position information and $M\in\mathbb{R}^{(T+1)\times(T+1)}$ is the attention mask. The total sequence length is increased to $T+1$ because of the use of special delimiter tokens ([B] or [E]). Next, MHA is used to establish position-context-image attention. as:
$Attn_{f}=Attn_{p}+dropout[MHA(Attn_{p},z_{l})],$ (11)
@@
$Attn_{p}=p+dropout[MHA(p,Attn_{cv},M)],$ (12)
where $p$ is the position queries. By updating the contextual semantic representation, each position token can gather the global information guided by the current input. CHA first helps the model capture dependencies between context and the visual information by allowing sequence interactions with the adaptive attention mask ($M$) provided by IPN. Then, the correlation between position and context is enhanced through context-image implicit guidance, thus getting rid of additional IR. The robustness of the model is improved because the uniform gradient flow allows IPN and CHA to interact with each other during training. Finally, the output logits $y^{\prime}$ are obtained as:
@@
where $S$ is the size of the character set used for training. Additional characters are associated with the [E] token (the end of the token sequence). In the optimization phase, the training loss is the average of the $K$ cross-entropy loss $\mathcal{L}_{ce}$ with attention mask, as:
@@ IV-A Dataset
This experiment uses both synthetic training datasets MJSynth (MJ) [14] and SynthText (ST) [13] as well as real data for training including nine real-world datasets (COCO-Text (COCO) [44], RCTW17 [37], Uber-Text (Uber) [58], ArT [7], LSVT [41], MLT19 [30], ReCTS [56], TextOCR [38] and OpenVINO [17]). We use IIIT 5k-word (IIIT5k) [28], CUTE80 (CUTE) [34], Street View Text (SVT) [46], SVT-Perspective (SVTP) [31], ICDAR 2013 (IC13) [16] and ICDAR 2015 (IC15) [15], COCO, Uber, ArT for testing. The dataset contains challenging texts such as multi-orientation, curved, blurry, low resolution, distorted, and occluded text.
@@
Synthetic datasets MJ consists of 8.9 million photorealistic text images composed of synthetic datasets through over 90k English dictionaries. The composition of MJ consists of three parts background, foreground, and optional shadow/border respectively. It uses 1400 different fonts. Additionally, the font word spacing, thickness, underlining, and other attributes of MJ are different. MJ also utilizes different background effects, border/shadow rendering, base shading, projection distortion, natural image blending, and noise. ST consists of 8 million word text images synthesized by blending text over natural images. It uses scene geometry, textures, and surface normals to naturally blend and distort text renderings on the surfaces of objects in the image. Similar to MJ, ST uses randomized fonts for its text. The text image is cropped from the natural image in which the synthesized text is embedded.
@@
Real-world datasets The COCO dataset has a total of 73,127 training and 9.8k test images containing non-text, legible, illegible, and occluded text images. Each image contains at least one instance of readable text. RCTW contains 12,263 annotated images of the large-scale Chinese field dataset. The Uber dataset has 127,920 training 60,000 test images collected from Bing Maps Streetside. It contains house numbers and vertical and rotated text on signboards. The ArT dataset was created to recognize arbitrarily shaped text containing 32,028 training images and 26,000 test images of arbitrarily shaped, perspective, rotated, or curved text. SVT is a large-scale street scene text dataset collected from streets in China, including 41,439 text images of arbitrarily shaped natural scenes. MLT19 was created to recognize multilingual text. It has a total of 56,727 training images consisting of seven languages: Arabic, Latin, Chinese, Japanese, Korean, Bengali and Hindi. The ReCTS dataset contains 26,432 irregular texts arranged in various layouts or written in unique fonts. TextOCR and OpenVINO are large datasets with diverse images containing 818,087 and 2,071,541 images, respectively. They have complex scenes with multiple objects and text of different resolutions, orientations, and qualities.
@@
The IIIT5k dataset is a collection of 5000 natural scenes and digitally born text images crawled from Google image search, containing 2,000 images for evaluation and 3,000 images for testing. It has low resolution, multiple font styles, light and dark transformations, and projection distortion. CUTE contains 288 cropped images for the curved text collection. The images were captured by digital cameras or collected from the Internet. SVT is a collection of street view text from Google Street View. It contains 257 images for evaluation and 647 images for testing. Further, the SVTP dataset is a more complex collection of images containing numerous perspective texts, totaling 645 images for testing. IC13 and IC15 were created for the ICDAR Robust Reading Contest. IC13 contains 848 images for training and 1,015 images for evaluation. IC15 contains 4,468 images for training and 2,077 images for evaluation. Many of the samples contain perspective and blurred text.
@@ IV-B Implementation Details
The entire training and testing is implemented in two RTX 3090 GPUs with 24GB of RAM. The training is performed with a batch size of 1024 for 31,815 iterations i.e. 10 epochs on a real dataset of 3,257,585 samples and 4 epochs on a synthetic dataset of 16.89M samples. Adam [18] optimizer is used along with the 1cycle [39] learning rate scheduler with an initial learning rate of 7e57e-5. We set a maximum label length of $T=25$ and used a character set of size $S=94$ containing mixed-case alphanumeric characters and punctuation. The image is augmented using a three-layer RandAugment [8] operation where sharpening is tuned to Gaussian Blur and Poisson Noise and resized to $128\times 32$ pixels. Each patch is set a size of $8\times 4$. For the model testing phase, we use a 36-character set i.e. containing numbers and 26 letters. We use word accuracy as the main evaluation metric i.e., the prediction is considered correct if and only if the characters match at all positions.
Remark The highlighted red and red lines in the visualization results indicate error prediction and lack of word, respectively. The best performance result is shown in bold font.
@@ IV-C Comparison with State-of-the-Arts
We compare HAAP with recent methods on 6 public benchmarks, and the quantitative results are listed in Table. II. We can see that HAAP outperforms the benchmark Parseq [4], realizing SOTA performance. The recognition performances in IC15 (incident scene text), SVTP (perspective scene text), and CUTE80 (curved text) are improved by $1.1\%$, $1.7\%$, and $1\%$ respectively. Our model achieves $99.3\%$ recognition accuracy on the IIIT5K (distorted, low-resolution scene text) dataset. This is an improvement of $2.6\%$ compared to the latest method, HVSI [9]. This verifies the excellent performance of HAAP on irregular text datasets, which contain a large number of low-quality images such as distorted and blurred images.
Besides validating the model on small-scale public benchmarks, we evaluate HAAP on three larger and more challenging recent benchmarks consisting of irregular text with different shapes, low-resolution, rotated, and occluded text. The results are shown in Table. III. It achieves $80.6\%$, $85.5\%$, and $85.8\%$ outperforms the previous methods in terms of accuracy, respectively. Representative visible results are shown in Fig. 5, which demonstrates that our method is sufficiently robust to occlusion and text direction variability. In detail, we find that HAAP and recent methods are effective in solving complex background problems caused by perspective shifts or uneven lighting. However, artistic lettering and occlusion can lead to errors in previous work. The independence of heterogeneous data leads to ambiguous judgments by the recognizer, which can be effectively avoided by HAAP. Furthermore, HAAP can protect against additional distortions and ambiguities. For example, although the second half of ”CORNER” in Fig. 5 is ambiguous, our model can still reason effectively. HAAP also recognizes that the ”R” in ”SHARP” in Fig. 5 relies on AR reasoning with adaptive permutations.
@@
TABLE IV: Comparison with other methods in terms of complexity and efficiency. AVG6AVG-6 represents the average accuracy in the six benchmarks.
@@
Ours 20.9 3.6 15.2 95.0
In addition, Fig. 6 and Table. IV shows the trade-off between accuracy and cost (parameters, FLOPS, and latency). For concise representation, we use $Avg$ to denote the average accuracy, which is computed as a weighted average based on the number of datasets. HAAP achieves the highest average word accuracy and exhibits competitive cost-quality in all three metrics. Compared to PARSeq, HAAP uses far fewer parameters and FLOPs. In terms of latency, HAAP achieves Avg6Avg-6 of 95.0% with 15.2 ms/image, which outperforms previous methods.
TABLE V: Ablation study of the IPN.Avg9Avg-9 represents the average accuracy in nine datasets. $P_{left2right}$ and $P_{right2left}$ represent the left-to-right and right-to-left permutation.
@@
IV-D Ablation study
To evaluate the effectiveness of the proposed HAAP with IPN and CHA, we conduct ablation studies on nine benchmark datasets. In these experiments, all models are trained using real datasets with consistent settings of their model hyperparameters.
The effectiveness of IPN We compare IPN with different sequence modeling principles (i.g, left-to-right, bi-directional, and PLM) to evaluate the effectiveness of adaptive sequence modeling. Table. V shows that the bidirectional model can achieve a performance gain of about $3.4\%$ on Avg9Avg-9 compared to the left-to-right model. Compared to the PLM model ($90.12\%$), the IPN model achieves a performance gain of $0.28\%$ in terms of Avg9Avg-9. To demonstrate the effectiveness of IPN in model training, we show the training and validation curves in five continuous training sessions in Fig. 7. Meanwhile, we use early stopping to determine the iteration step of model convergence by validating that the increase of Avg9Avg-9 is less than $0.3\%$ in 2000 steps. The result is depicted in Tab. VI. Specifically, as shown in Fig. 7, the IPN model can achieve more stable training and convergence than the PLM model in multi-training. As depicted in Tab. VI, the Avg9Avg-9 standard deviation (Std Dev) of the IPN model is $0.33$, which is much less than the PLM model ($1.45$). With IPN, adaptive sequence modeling can help the model avoid train fit oscillation and achieve convergence advantage with less training data and time costs than PLM. The qualitative results in Fig. 8 show that the IPN model can effectively predict some complex cases influenced by background.
(a) the model with PLM
(b) the model with IPN
Figure 7: Comparison of Loss (Left) and Valuation (Right) curves in five continuous training sessions.
TABLE VI: Training schedule comparison in five continuous training sessions. Std Dev represents standard deviation.
Training session Step of convergence Avg-9 (%) Sample count (M) Training hours
PLM
Session 1 15318 90.12 15.69 5.23
Session 2 19591 88.89 20.06 5.79
Session 3 18500 86.83 18.94 5.64
Session 4 20091 87.62 20.57 5.86
Session 5 999 0 10.23 2.6
Mean 18375 88.37 18.82 5.63
Std Dev 2143.53 1.45 2.19 0.28
IPN
Session 1 11636 90.37 11.92 4.67
Session 2 12545 89.67 12.85 4.79
Session 3 12545 89.76 12.85 4.79
Session 4 13727 89.78 14.06 4.91
Session 5 12545 90.31 12.85 4.79
Mean 12599 89.98 12.91 4.79
Std Dev 743.05 0.33 0.76 0.08